NaIO3 + 5NaI + 2H3C6H5O7 3I2 + 3H2O + 2Na3C6H5O7



=? g/ml


1
Expert's answer
2022-01-30T07:57:17-0500

This is a redox reaction

"5\>I^-\to\>5\>I+5e^-"

"I^{5+}+5e^-\to\>I"


Molar mass of citric acid "=192.124g\/mol"

Molar mass of Iodine "=126.9g\/mol"



If 2.5 of iodine is deposited by using 25ml of citric acid


Then concentration of citric acid is;

Moles of iodine deposited "=\frac{2.5}{126.9}=0.01970"


Moles of citric acid "=\frac{2}{3}×0.01970=0.01313"


Concentration of citric acid"=0.01313×\frac{1000}{25}×192.124"

"=100.9g\/litre"

"=0.1009g\/ml"